OpenStudy (anonymous):

Determine the slope and the y-intercept for the linear equation. -X+3Y-12=0

OpenStudy (anonymous):

Isolate y = ...

OpenStudy (anonymous):

I calculated slope m=1/3 and I don't know how to find y-intercept

OpenStudy (anonymous):

Could you show your work HERE?

OpenStudy (anonymous):

-x+3y-12=0 3y=x+12 y=1/3x+12/3 y=1/3x+4 According to formula y= mx+b m=1/3

OpenStudy (anonymous):

y= 1/3x + 4 So m=1/3, y = +4

OpenStudy (anonymous):

b is y intercept!

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Learn the terms, you'll see algebra is easy to you :)